OpenStudy (anonymous):

CRAFTS Madelyn cut a 60-inch pipe cleaner into two unequal pieces, and then she used each piece to make a square. The sum of the areas of the squares was 117 square inches. Let x= the length of one piece. Write and solve an equation to represent the situation and find the lengths of the two original pieces.

Directrix (directrix):

Here is where we left off. Read the info at the link first. (x/4)² + [(60 -x)/4]² = 117 Solve for x.

OpenStudy (anonymous):

does that equal 2x^2 -120x+243=0 ? and from there solve for x with quadratic formula ?

Directrix (directrix):

I think I have: x^2 - 60x +1800 = 0 Let's check our products.

Directrix (directrix):

Mine is wrong.

Directrix (directrix):

I think your product is also incorrect.

Directrix (directrix):

Let us multiply again.

OpenStudy (anonymous):

i got x^2/16 + x^2-120x+360/16 = 117

Directrix (directrix):

(x/4)² + [(60 -x)/4]² = 117 x^2/16 + (x^2 -120x + 360) /16 = 117

Directrix (directrix):

Let's multiply this one more time: [(60 -x)/4]² = [(60 -x)/4] * [(60 -x)/4] = (60 - x)^2 / 16

Directrix (directrix):

(60 - x)^2 / 16 = This part is: (60 - x)^2 = 3600 - 120x + x^2

Directrix (directrix):

x^2/16 + (x^2 -120x + 3600) /16 = 117 Multiply through by 16

Directrix (directrix):

Eventually, it factors into (x - 24)*(x-36) = 0

OpenStudy (anonymous):

isn't 16 a common denominator because it's adding fractions?

Directrix (directrix):

Yes, it is.

Directrix (directrix):

it factors into (x - 24)*(x-36) = 0 so x = 24 or x = 36. Which ones of those makes sense in the context of the problem?

OpenStudy (anonymous):

I don't get how you got to that stage .. from x^2/16 + (x^2 -120x + 3600) /16 = 117 doesn't it reduce to x^2+ X^-120x+3600 = 117 so you subtract 117 to 3600?

Directrix (directrix):

Did we agree up to this point? x^2/16 + (x^2 -120x + 3600) /16 = 117

OpenStudy (anonymous):

yeess

Directrix (directrix):

So far, so good up until that point.

OpenStudy (anonymous):

or is it x^2-120x+36-/16 = 117 so you multiply 16 with 117 to get 2x^2-120x-1512=0 ?

OpenStudy (callisto):

Check the constant term. It is not correct.

OpenStudy (anonymous):

from that i got x=42 im so confuseeed

OpenStudy (callisto):

\[\frac{ x^2}{16} + \frac{(x^2 -120x + 3600)}{16} = 117\]\[\frac{ x^2+(x^2 -120x + 3600)}{16} = 117\]Multiply both sides by 16 \[16\times \frac{ x^2+(x^2 -120x + 3600)}{16} = 16\times 117\]

OpenStudy (anonymous):

oooh i understand 9x-34) (x-36) thank you (: my calc. was wrong

Directrix (directrix):

x^2/16 + (x^2 -120x + 3600) /16 = 117 ------------------------------- Multiply by 16. x^2 + (x^2 -120x + 3600) = 16*117 2 x^2 - 120 x + 3600 = 1872 @jenna_kwak Did you get this. I lost a denominator along the way and then multiplied incorrectly.

OpenStudy (anonymous):

i got 2x^2-120x+3600=1872 subtracted and got 2x^2-120x+1728 factored 2(x^2-60x+864) and got (x-24) and (x-36)

Directrix (directrix):

2 x^2 - 120 x + 3600 = 1872 2 x^2 - 120 x + 1728 = 0 x^2 - 60 x + 864 = 0 (x-24)* (x-36) = 0 x = 24 and x = 36. Such a good feeling when it finally all works out.

OpenStudy (anonymous):

yes (: the heart of math thank you gus so much im in precal and my algebra1 sister didnt get this and i was struggling for like a hour -____- thank you !

Directrix (directrix):

The struggle is how we get tough and good at solving these things. You never gave up. We'll give @Callisto a nod for catching that arithmetic error :).

Directrix (directrix):

I enjoyed the process. Now, explain it to your sister.

OpenStudy (anonymous):

i am thank you (: